The time, in minutes, it takes Helen to get to work varies inversely as speed. If she can get to work in 25 minutes at a speed of 48 mph, which of the following equations best models the time, , it will take her to get to work at a speed of mph?

Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)



If you are going to ask a question that contains "...which of the following...", then you have to list the answers you were given. Either that, or write to the Psychic Hot Line instead of Algebra.com.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Alan3354(69443)

The time, in minutes, it takes Helen to get to work varies inversely as speed. If she can get to work in 25 minutes at a speed of 48 mph, which of the following equations best models the time, , it will take her to get to work at a speed of mph?
---------------
d = r*t = (25/60)*48 = 20 miles
-----
t = d/r
t = 20/r, t in hours, r in mi/hr